# VLM-E2E: Enhancing End-to-End Autonomous Driving with Multimodal Driver Attention Fusion

### v1

@@ III-D Attention Guided Future Planning
The primary objective of the proposed motion planner is to generate trajectories that ensure safety, comfort, and efficient progress toward the goal. To achieve this, we employ a motion planner that generates a set of kinematically feasible trajectories, each of which is evaluated using a learned scoring function, inspired by [63, 64, 65, 43].
Our scoring function incorporates a probabilistic dynamic occupancy field, which is crucial for encoding the safety of the potential maneuvers. This field encourages cautious driving behaviors by penalizing trajectories that enter occupied regions or get too close to these regions, thus maintaining a safe distance from surrounding obstacles. Additionally, we utilize the probabilistic layers from our online map to inform the scoring function. These layers provide important information, ensuring that the self-driving vehicle (SDV) remains within the drivable area, stays close to the center of the lane, and moves in the correct direction. Particularly in regions of uncertainty, where occupancy and road structure are less predictable, the planner takes extra care to drive cautiously. Moreover, the planner ensures that the vehicle progresses toward the goal specified by the input high-level command, whether it is to continue forward, make a turn, or navigate other maneuvers.
The planner evaluates all sampled trajectories in parallel. Each trajectory $\tau$ is assessed based on the scoring function $f$, which considers several input factors, including the map $\mathcal{M}$, occupancy $\mathcal{O}$, and the motion $\mathcal{V}$. The trajectory selection process is formulated as:
[…]
where $\tau^{*}$ denotes the optimal trajectory, $f(\tau,\mathcal{M},\mathcal{O},\mathcal{V},w)$ is the learned scoring function, $w$ are the learnable parameters of the model.
The scoring function evaluates each trajectory concerning multiple criteria, such as the safety of the maneuver avoiding obstacles, the comfort of the ride such as maintaining smooth motion, and progress toward the goal, as guided by the high-level command. By combining these factors, the motion planner efficiently selects the trajectory that best satisfies all safety, comfort, and progress criteria, ensuring the SDV navigates complex environments in a manner that is both effective and cautious.
The output of the motion planner is a sequence of vehicle states, which defines the desired motion of the SDV within the planning horizon. In each iteration of the planning process, a set of candidate trajectories is generated and evaluated using the cost function described in (11). The output of the motion planner is a sequence of vehicle states, which defines the desired motion of the SDV within the planning horizon. The trajectory with the minimum cost is then selected for execution.
@@
To ensure the planned trajectory adheres to driver attention on traffic regulations and route, we utilize a temporal refinement module that dynamically integrates traffic regulations. Leveraging front-view camera features $e_{front}$ from the encoder, we initialize a GRU-based refinement network to iteratively adjust the initially selected trajectory. The front-view features explicitly encode traffic regulations semantics, enabling the model to halt at red lights or proceed through green signals. The recurrent architecture ensures smooth transitions between trajectory points, mitigating abrupt steering or acceleration changes.
IV Experimental Settings
IV-A Dataset
@@
We evaluate our method on the nuScenes dataset [68], a large-scale autonomous driving benchmark comprising 1,000 diverse driving scenes, each spanning 20 seconds with annotations provided at 2 Hz. The dataset features a 360° multi-camera rig composed of six synchronized cameras (front, front-left, front-right, back, back-left, back-right) with minimal field-of-view overlap. Precise camera intrinsic and extrinsic are provided for each frame to ensure accurate spatial alignment.
@@
The BEV occupancy labels $\{y_{t+1},\cdots,y_{t+l}\}$ are generated by projecting 3D bounding boxes of dynamic agents onto the BEV plane, creating a spatiotemporal occupancy grid. All labels are transformed into the ego vehicle’s reference frame using GT future ego-motion, ensuring temporal consistency across frames.
IV-B Metrics
The perception performance is assessed using the Intersection over Union (IoU), which quantifies the overlap between predicted and GT object bounding boxes. This metric is commonly used to evaluate the accuracy of object detection and tracking in autonomous driving systems.
$IoU=\frac{A\cap B}{A\cup B}$ (12)
where $A$ represents the predicted segmentation, $B$ represents the ground truth segmentation.
For prediction evaluation, we employ three key metrics. Panoptic Quality (PQ) evaluates the overall quality of both semantic segmentation and instance detection, accounting for both the accuracy of object classification and the correctness of instance segmentation. Recognition Quality (RQ) measures the ability of the model to correctly recognize and classify objects within the scene. Segmentation Quality (SQ) focuses on the accuracy of the predicted segmentation masks, comparing them with the GT to evaluate the precision of the object segmentation.
$PQ=\underbrace{\frac{\sum_{(p,g)\in TP}IoU(p,g)}{|TP|}}_{\text{Segmentation Quality (SQ)}}\times\underbrace{\frac{|TP|}{|TP|+\frac{1}{2}|FP|+\frac{1}{2}|FN|}}_{\text{Recognition Quality (RQ)}}$ (13)
where $TP$, $FP$, and $FN$ refer to true positives, false positives, and false negatives, respectively. $p,g$ are the predicted and GT instances.
For evaluating the planning performance, we consider two primary metrics. L2 Distance measures the Euclidean distance between the SDV’s planned trajectory and real human-driven trajectories. A lower L2 distance indicates that the model is better able to replicate human-like driving behavior, which is important for ensuring natural and comfortable motion. Collision Rate quantifies the percentage of time steps during a trajectory in which the SDV’s predicted path collides with the ground truth bounding box of other agents. It provides an indication of the safety of the planner’s decisions.
$CR=\frac{1}{T}\sum_{t=1}^{T}\mathrm{1}(\tau_{t}^{*}\cap o_{t}\neq\emptyset)$ (14)
where $\tau^{*}$ is the planned trajectory and $o$ is the other agents’ ground-truth occupancy.
$L2=\frac{1}{T}\sum_{t=1}^{T}||\tau_{t}^{*}-\tau_{t}^{human}||_{2}$ (15)
where |||||$\cdot$||| | ⋅ | | is the Euclidean distance between the SDV’s planned trajectory at time $t$ and the corresponding human trajectory at the same time.
IV-C Implementation Details
Our model utilizes a temporal context of 1.0 seconds of past information to predict the future trajectory over a 2.0-second horizon. In the nuScenes dataset, this corresponds to 3 frames of past context and 4 frames into the future, operating at a frequency of 2 Hz.
At each past timestep, the model processes 6 camera images, each with a resolution of 224 × 480 pixels. The BEV spatial is $100m\times 100m$ area at a pixel resolution of 50cm in both the x and y directions. This results in a BEV video with spatial dimensions of $200\times 200$ pixels.
Training is performed using the Adam optimizer with a constant learning rate of $2.0\times 10^{-3}$. The model is trained for 20 epochs with a batch size of 6, distributed across 4 Tesla A6000 GPUs. To optimize memory usage and accelerate computation, mixed precision training is employed. Additionally, both our model and ST-P3 are trained without depth guidance, ensuring a fair comparison and highlighting the effectiveness of our approach in leveraging semantic and attentional cues for improved performance.
V Results
V-A Quantitative Results
V-A1 Perception
Table I presents the perception performance of various methods across four key categories: Drivable Area, Lane, Vehicle, and Pedestrian. Our proposed VLM-E2E model demonstrates significant improvements over existing approaches, achieving best results in three out of four categories. Specifically, VLM-E2E outperforms ST-P3 in lane detection with a $2.24\%$ relative improvement, vehicle detection with an $0.75\%$ increase, and Pedestrian detection with a $24.40\%$ boost on the nuScenes validation set. While IVMP achieves the highest score in drivable area detection, VLM-E2E closely follows with a score of 74.69, demonstrating competitive performance. These results highlight the effectiveness of our end-to-end approach in enhancing perception accuracy, particularly in critical tasks such as lane and vehicle detection, which are essential for safe and reliable autonomous driving.
@@ V-A1 Perception
V-A2 Prediction
Table II presents the prediction performance of various methods on the task of semantic and instance segmentations in BEV for a future horizon of 2.0 seconds. We evaluate the methods using IoU, PQ, SQ, and RQ. Our proposed method, VLM-E2E, achieves best performance across IoU, PQ, and RQ. Specifically, VLM-E2E attains an IoU of 38.54, representing a $4.47\%$ improvement over ST-P3. In terms of PQ, VLM-E2E achieves a PQ of 29.83, demonstrating superior instance segmentation capabilities compared to ST-P3 and FIERY. These results highlight the effectiveness of VLM-E2E in capturing both semantic and instance-level information in dynamic driving scenarios.
@@ V-A2 Prediction
Method IoU PQ SQ RQ
@@
V-A3 Planning
The results presented in Table III illustrate the performance of various methods in planning tasks, specifically focusing on L2 displacement error and collision rate across 1s, 2s, and 3s time horizons. Notably, VLM-E2E exhibits significant improvements, particularly in longer-term planning scenarios. For the L2 displacement error, VLM-E2E achieves the best performance at the 3s horizon with a score of 2.68, outperforming all other methods. While Vanilla demonstrates superior performance at the 1s and 2s horizons, VLM-E2E’s notable reduction in long-term prediction errors is critical for ensuring reliable autonomous driving.
In terms of collision rates, VLM-E2E achieves state-of-the-art performance, recording the lowest rates of 0.60% at the 2-second horizon and 1.17% at the 3-second horizon, surpassing all other methods. Although ST-P3 achieves a slightly lower collision rate at the 1-second horizon, the consistent improvements demonstrated by VLM-E2E over longer time underscore its robustness in minimizing collision risks during extended planning periods. Overall, VLM-E2E demonstrates substantial advancements in long-term planning accuracy and safety, particularly in reducing both L2 displacement errors and collision rates at the 2s and 3s horizons. These results highlight VLM-E2E’s potential to enhance planning systems in autonomous driving, especially in scenarios that demand reliable long-term predictions and effective collision avoidance.
TABLE III: Planning results. We report the L2 (m) and CR (%) across 1s, 2s, 3s.
Method L2 (m) CR (%)
1s 2s 3s 1s 2s 3s
Vanilla [72] 0.50 1.25 2.80 0.68 0.98 2.76
NMP [73] 0.61 1.44 3.18 0.66 0.90 2.34
Freespace [11] 0.56 1.27 3.08 0.65 0.86 1.64
ST-P3 [43] 1.33 2.11 2.90 0.23 0.62 1.27
VLM-E2E 1.22 1.94 2.68 0.26 0.60 1.17
@@ V-A3 Planning
Figure 3: Qualitative analysis on prediction. (a) shows the multi-view input images. (b) shows the heatmap (blue to red) which illustrates the probability distribution of instance centers within the scene, with warmer colors indicating higher confidence regions. (c) represents the vehicles segmentation which effectively distinguishes individual instances in the complex traffic scenario. (d) reveals the directional vectors pointing towards the corresponding instance centers for each pixel, demonstrating the model’s understanding of spatial relationships. (e) exhibits consistency within each instance, reflecting the characteristic rigid-body motion of vehicles.
@@
V-B Qualitative Analysis
Fig. 3 demonstrates the generated outputs, including instance segmentation, instance center, instance offset, and future flow. Fig. 3(b) features a heatmap highlighting detected objects, while Fig. 3(c) displays the instance segmentation results, where each segment is color-coded to represent different objects. The offset is a vector pointing to the center of the instance in Fig. 3(d). The future flow Fig. 3(e) is a displacement vector field of the dynamic agents. These visualizations enhance the understanding of spatial relationships and the distribution of elements within the environment, underscoring the model’s capability to accurately perceive and segment critical features essential for autonomous driving applications.
VI Conclusion
In this paper, we propose VLM-E2E, a novel end-to-end autonomous driving framework that leverages VLMs to enhance semantic understanding of driver attention. Our approach is motivated by the need to address key limitations in existing systems, such as modality imbalance in multi-sensor fusion, insufficient utilization of high-level semantic context, and the lack of interpretability in trajectory planning. To this end, we introduce a BEV-Text learnable weighted fusion strategy to dynamically balance geometric and semantic features, a spatiotemporal module to ensure temporal coherence in dynamic scenes and a probabilistic future prediction module with attention guiede trajectory refinement. These components collectively enable our framework to achieve robust and interpretable performance across perception, prediction, and planning tasks. Future work will focus on extending the framework to incorporate VLMs and E2E into a unified framework and utilize the lidar and radar modalities in our framework to generalize our model in long tail scenarios.

### v2

@@ III-D Attention Guided Future Planning
The primary objective of the proposed motion planning is to generate trajectories that ensure safety, comfort, and efficient progress toward the goal. To achieve this, we employ a motion planner. The planner generates a set of kinematically feasible trajectories, each of which is evaluated using a learned scoring function, inspired by [31, 32, 33, 16].
@@
To ensure the planned trajectory adheres to driver attention on traffic regulations and routes, we utilize a temporal refinement module that dynamically integrates traffic regulations. Leveraging front-view camera features $e_{front}$ from the encoder, we initialize a GRU-based refinement network to iteratively adjust the initially selected trajectory. The front-view features explicitly encode traffic regulations semantics, enabling the model to halt at red lights or proceed through green signals. The recurrent architecture ensures smooth transitions between trajectory points, mitigating abrupt steering or acceleration changes.
IV Experiments
IV-A Experimental Settings
@@
Dataset. We evaluate our method on the nuScenes dataset [36], a large-scale autonomous driving benchmark comprising 1,000 diverse driving scenes, each spanning 20 seconds with annotations provided at 2 Hz. The dataset features a 360° multi-camera rig comprising six synchronized cameras with minimal field-of-view overlap. Precise camera intrinsic and extrinsic are provided for each frame to ensure accurate spatial alignment.
@@
All labels are transformed into the ego vehicle’s reference frame using GT future ego-motion, ensuring temporal consistency across frames. Besides, maneuvering is used to correct annotations from VLMs.
Implementation Details. Our framework processes 1 second of historical sensor data, equivalent to 3 frames at 2Hz, to predict trajectories over a 3 s horizon spanning 6 future frames. Inputs include synchronized multi-camera RGB streams with $224\times 480$ pixel resolution and BEV grid maps covering a 100m×100m area at 0.5m per pixel resolution. Training employs the Adam optimizer with a learning rate of $2.0\times 10^{-3}$ for 20 epochs, using mixed precision on 4 NVIDIA A6000 GPUs and a batch size of 6.
@@
Method IoU $\uparrow$ PQ $\uparrow$ SQ $\uparrow$ RQ $\uparrow$
@@
TABLE III: Planning results. We report the L2 (m) and Collision Rate CR (%) across 1s, 2s, 3s.
Method L2 (m) $\downarrow$ CR (%) $\downarrow$
1s 2s 3s Avg. 1s 2s 3s Avg.
Vanilla [43] 0.50 1.25 2.80 1.52 0.68 0.98 2.76 1.47
NMP [44] 0.61 1.44 3.18 1.74 0.66 0.90 2.34 1.30
Freespace [6] 0.56 1.27 3.08 1.64 0.65 0.86 1.64 1.05
ST-P3 [16] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71
UniAD [2] 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
VAD [1] 0.54 1.15 1.98 1.22 0.04 0.39 1.17 0.53
GenAD [45] 0.28 0.49 0.78 0.52 0.08 0.14 0.34 0.19
Senna [27] 0.37 0.54 0.86 0.59 0.09 0.12 0.33 0.18
VLM-E2E 0.28 0.50 0.80 0.53 0.01 0.06 0.20 0.09
IV-B Quantitative Results
Perception. Table I presents the perception performance of various methods across four key categories: Drivable Area, Lane, Vehicle, and Pedestrian. Our proposed VLM-E2E model demonstrates significant improvements over existing approaches, achieving the best results in three out of four categories. Specifically, VLM-E2E outperforms ST-P3 in lane detection with a $2.24\%$ relative improvement, vehicle detection with an $0.75\%$ increase, and Pedestrian detection with a $24.40\%$ boost on the nuScenes validation set. While IVMP achieves the highest score in drivable area detection, VLM-E2E closely follows with a score of 74.69, demonstrating preserved geometric reasoning in BEV space. These results demonstrate that integrating driver-attentional features enhances critical perception tasks without compromising foundational scene understanding.
@@ IV-B Quantitative Results
Figure 3: Qualitative analysis on prediction. (a) shows the multi-view input images. (b) shows the heatmap (blue to red) which illustrates the probability distribution of instance centers within the scene, with warmer colors indicating higher confidence regions. (c) represents the vehicles segmentation, which effectively distinguishes individual instances in the complex traffic scenario. (d) reveals the directional vectors pointing towards the corresponding instance centers for each pixel, demonstrating the model’s understanding of spatial relationships. (e) exhibits consistency within each instance, reflecting the characteristic rigid-body motion of vehicles.
Prediction. Table II presents the prediction performance of various methods on the task of semantic and instance segmentations in BEV for a future horizon of 2.0 seconds. We evaluate the methods using IoU, PQ, SQ, and RQ. VLM-E2E achieves the best performance across IoU, Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). Specifically, VLM-E2E attains an IoU of 38.54, representing a $4.47\%$ improvement over ST-P3. In terms of PQ, VLM-E2E achieves a PQ of 29.83, demonstrating superior instance segmentation capabilities compared to ST-P3 and FIERY. These results highlight the effectiveness of VLM-E2E in capturing both semantic and instance-level information in dynamic driving scenarios.
Planning. As shown in Table III, our VLM-E2E achieves the best overall performance among all methods. It achieves the lowest average L2 error, 0.53 m, and substantially reduces the collision rate to 0.09%. Compared to previous state-of-the-art approaches, such as GenAD, which achieves 0.52 meters and 0.19 percent, and Senna, which reaches 0.59 m and 0.18%, VLM-E2E delivers notable improvements in both trajectory accuracy and safety. Moreover, the gains over traditional baselines like Vanilla and NMP are even more significant. These results clearly demonstrate the effectiveness of incorporating VLM-guided attention into end-to-end autonomous driving frameworks.
TABLE IV: Ablations of text encoders. We report the semantic segmentation IoU (%) in BEV.
Method Drivable Area Lane Vehicle Pedestrian
Bert 73.44 38.19 38.53 15.90
CLIP 74.69 39.33 39.08 17.49
TABLE V: Ablations of input text views. We report the semantic segmentation IoU (%) in BEV.
Text View Drivable Area Lane Vehicle Pedestrian
All View 73.50 37.98 38.79 17.51
Front View 74.69 39.33 39.08 17.49
IV-C Qualitative Analysis
Fig. 3 demonstrates the generated outputs, including instance segmentation, instance center, instance offset, and future flow. Fig. 3(b) features a heatmap highlighting detected objects, while Fig. 3(c) displays the instance segmentation results, where each segment is color-coded to represent different objects. The offset is a vector pointing to the center of the instance in Fig. 3(d). The future flow Fig. 3(e) is a displacement vector field of the dynamic agents. These visualizations enhance the understanding of spatial relationships and the distribution of elements within the environment, underscoring the model’s capability to accurately perceive and segment critical features essential for autonomous driving applications
@@ IV-C Qualitative Analysis
IV-D Ablation Study
Text Encoder. Table IV compares the performance of text encoders, Bert [46] and CLIP, across four detection categories. Bert shows a consistent performance decline, with decreases of 1.67%, 2.90%, 1.41%, and 9.09%, respectively, highlighting its limitations in perception tasks. In contrast, CLIP demonstrates superior capability in encoding textual annotations, achieving higher accuracy across all categories. The results validate that CLIP’s joint embedding space effectively mitigates modality gaps between text annotations and BEV representations.
Input Text View. Table V evaluates the impact of text view selection on BEV semantic segmentation performance. When transitioning from all view to front view text inputs, we observe performance improvements of 1.62% for drivable area, 3.55% for lane, and 0.75% for vehicle, respectively. This improvement primarily stems from eliminating interference caused by redundant textual context in all views, where excessive multi-view descriptions introduce semantic noise that distracts the model from task-critical features. The results validate our design choice to adopt the front view as the default configuration, as it optimally balances semantic specificity and noise suppression for scene understanding tasks.
V Conclusion
We propose VLM-E2E, an E2E autonomous driving framework that leverages VLMs to enhance driver-attentional semantic understanding. We introduce a BEV-Text learnable weighted fusion strategy balancing geometric-semantic features to address modality imbalance and semantic underutilization in existing systems, along with a spatiotemporal coherence mechanism for temporal consistency, and an attention-guided trajectory refinement for probabilistic prediction. Future work will extend integration with advanced E2E architectures for improved generalization.